OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!! Bonnie has one alloy containing 23% aluminum and another containing 45% aluminum. How many pounds of each alloy must he use to make 58 pounds of a third alloy containing 42% aluminum?

OpenStudy (anonymous):

42% of 58= \[.42\times 58=24.36\] so that is how much you aluminum you want to end up with. put x = amount of first compound which will have 23% of x = \[.23x\] aluminum. the remainder is 58-x which will have \[.45(58-x)\] aluminum for a total of \[.23x+.45(58-x)\] aluminum. set this total to equal 24.36 and get the equation \[.23x+.45(58-x)=24.36\]

OpenStudy (anonymous):

the work now is to solve this equation for x. i would multiply by 100 to clear the decimals and rewrite it as \[23x+45(58-x)=2436\]

OpenStudy (anonymous):

\[23x+2610-45x=2436\] \[-22x+2610=2436\] \[-22x=2436-2610\] \[-22x=-174\] \[x=\frac{174}{22}\] \[x=7.9\] rounded

OpenStudy (anonymous):

so about 7.9 pounds of the 23% aluminum and about 50.1 of the 45% aluminum. hope steps are clear